OpenStudy (lexy148):

Gram Formula Mass Zn(C2H3O2)2 x 2H2O i need help on how to add the #'s inside the ( )

OpenStudy (***[isuru]***):

do u want to find the molar mass inside ( ) ?

OpenStudy (lexy148):

the whole problem please

OpenStudy (***[isuru]***):

ok.... to calculate the gram formula mass of (C2H3O2)2 [ which u say u need help of ] first find the no.of atoms in that part there r C = 2 x 2 = 4 H = 2 x 3 = 6 O = 2 x 2 = 4  now find it in the usual way the relative atomic mass of Zn - 65 C - 12 H - 1 O - 16 so the Gram formula mass of Zn(C2H3O2)2 = 1zn + 4C + 6H + 4O = 65 + (4x12) + (6x1) + (4x16) = 65 + 48 + 6 + 64 = 183 g but u have to find the G F M of Zn(C2H3O2)2 x H2O which mean u have to add the G F M of H2O to 183 g G F M of H2O = 2H + O = (2x1) + 16 = 18 then the G F M of Zn(C2H3O2)2 x H2O = 183 + 18 = 201 g hope this will help ya!